Suppose a market has two firms that sell identical products. These firms face an inverse market...

  1. Suppose a market has two firms that sell identical products. These firms face an inverse market demand function of P=120 – Q. Firm 1 has a constant MC=20. Firm 2’s marginal cost is MC=30.
    1. Find the Cournot equilibrium price, quantities, and profits for each firm.
    2. If these firms were able to perfectly collude, what would be the monopoly equilibrium?
Answer #1

P-120-0 = 12000,-02 MC = 20 Me=30 TIF 1200.-( 20. 0₂ - 200, - 120-20, -0₂ - 2020 (0,=1000020 DO TIL=12002 - 0 h - 0 2 - 304 a